THE VIBRATION-ROTATION EMISSION SPECTRA OF GASEOUS CdH_{2} AND CdD_{2}

\underline{SHANSHAN YU}, ALIREZA SHAYESTEH, and PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

The vibration-rotation emission spectra of CdH_{2} and CdD_{2} molecules have been recorded at high resolution using a Fourier transform spectrometer. The molecules were generated in a furnace-discharge emission source by reaction of cadmium vapor with molecular hydrogen or deuterium. The fundamental bands for the antisymmetric stretching mode (ν_{3}) of CdH_{2} and CdD_{2} were detected at about 1771.5 cm^{−1} and 1278.3 cm^{−1}, respectively. In addition, the 002(Σ_{g}^{+}) – 001(Σ_{u}^{+}) and 011(Π_{g}) – 010(Π_{u}) hot bands were observed for CdH_{2}. Spectroscopic constants were determined for each of the twelve observed isotopologues: ^{110}CdH_{2}, ^{111}CdH_{2}, ^{112}CdH_{2}, ^{113}CdH_{2}, ^{114}CdH_{2}, ^{116}CdH_{2}, ^{110}CdD_{2}, ^{111}CdD_{2}, ^{112}CdD_{2}, 113CdD_{2}, ^{114}CdD_{2}, ^{116}CdD_{2}. The average Cd–H and Cd–D bond distances (r_{0}) were determined to be 1.683028(10) Å and 1.679161(16) Å, respectively.